OpenStudy (anonymous):

PLEASE HELP!!! The graph of a system of equations will intersect at more than 1 point. a. Always b. Sometimes c. Never

OpenStudy (debbieg):

It depends on the equations. If you are dealing with linear equations, the answer is different than if you allow non-linear equations. As stated, there is no restriction that the equations be linear. Consider the system: y=x^2 y=x+5 Think about what the graphs would look like.

OpenStudy (anonymous):

As debbie said it really depends on the functions in question. Two parallel lines will ALWAYS intersect if they are the same line, but will never intersect if they're not the same line, two arbitrary function would connect either once or more times, but again it depends on the functions.